OpenStudy (anonymous):

convert each equation to standard form by completing the square on x and y. Then graph the ellipse and give the location of its foci. 4x^2+25y^2-24x+100y+36=0 i need the steps

OpenStudy (anonymous):

because i did the problem but i think i messed up

OpenStudy (anonymous):

\[4x^2-24x+25y^2-100y=-36\] \[4(x^2-8x)+25(y^2-4y)=-36\] \[4(x-4)^2+25(y-2)^2=-36+4\times 16+25\times 4\] \[4(x-4)^2+25(y-2)^2=128\]

OpenStudy (anonymous):

4*8=32

OpenStudy (anonymous):

\[4(x-3)^2+25(y+2)^2\] right hand side should be \[-36+4\times 9+25\times 4=100\]

OpenStudy (anonymous):

divide by 100 get \[\frac{(x-3)^2}{25}+\frac{(y+2)^2}{4}=1\]

OpenStudy (anonymous):

wow did i mess up sorry. but last step is right

OpenStudy (anonymous):

\[4x^2-24x+25y^2+100y=-36\] \[4(x^2-6x)+25(y^2+4y)=-36\] \[4(x-3)^2+25(y+2)^2=-36+4\times 9+25\times 4\] \[4(x-3)^2+25(y+2)^2=100\]